# Dynamic correlations of frustrated quantum spins from high-temperature expansion

### v3

@@ Abstract
Here we narrow this methodological gap by introducing a dynamic extension of the high-temperature expansion (HTE). Conventional HTE targets equal-time correlations and thermodynamic observables via graph-based expansions in $J/T$ and is developed for more than five decades [20, 21]. It benefits from a formulation in the thermodynamic limit and is oblivious to frustration, entanglement, dimensionality and spin-length [22]. Its main limitation is the model-dependent temperature range $T\gtrsim O(J/4)$. Our dynamic HTE (Dyn-HTE) shares the same benefits but applies to frequency moments of the real-frequency dynamic susceptibility from which we reconstruct the DSF [23, 24, 25]. Although these moments are considerably more complex than equal-time correlators the high expansion orders of conventional HTE are maintained. This allows for the computation of four to seven frequency moments, depending on $T/J$, and goes beyond numerical linked-cluster expansions (NLCE) [26, 27] which computed the lowest two moments (Gauss approximation) for some particular S=1/2S\!=\!1/2 models [28].
@@
$M_{ii^{\prime},n}\!\!=\!i^{n}\partial_{t}^{n}\!\!\left\langle S_{i}^{z}(t)S_{i^{\prime}}^{z}\right\rangle\!|_{t=0}\!=\!\!\left\langle[...[S_{i}^{z},\!H],\!H],...,\!H]S_{i^{\prime}}^{z}\right\rangle\!.\!\!$ (6)
@@
Dyn-HTE for moments.—We now present our main conceptual result relevant for finite $T$: The HTE of moments $m_{\mathbf{k},2r}$ is encoded in the Dyn-HTE of the imaginary-frequency Matsubara spin correlator. This is significant as the latter treats $\rho$ and imaginary-time evolution $S_{i}^{z}(\tau)=e^{H\tau}S_{i}^{z}e^{-H\tau}$ on equal footing [33] and allows for an efficient evaluation of exact HTE coefficients, see our companion article [30]. The Matsubara correlator at frequency $\nu_{m}=2\pi Tm$ ($m\in\mathbb{Z}$) and its HTE to order $n_{max}$ in $x=J/T$ read
@@
Here $\mathcal{T}$ enforces $\tau$-ordering, $\Delta_{2\pi m}\equiv(1-\delta_{0,m})/(2\pi m)$ and $r_{max}=\left\lfloor n_{max}/2\right\rfloor$. The polynomials $p_{ii^{\prime}}^{(2r)}(x)$ will be linked to the HTE of the moments momentarily. They are of degree $n_{max}-2r$ with rational coefficients. Our open-source implementation [29] for $n_{max}=12$ provides $p_{ii^{\prime}}^{(2r)}(x)$ for all site-pairs $ii^{\prime}$ on arbitrary lattices $\mathcal{L}$. This flexibility hinges on the pre-evaluation of the Dyn-HTE (8) for abstract graphs which can then be embedded in any $\mathcal{L}$ like in conventional HTE [21]. The efficient evaluation of all $\simeq 1.6\times 10^{6}$ graphs with up to $n_{max}$ edges $(\sim\!J)$ [currently for $S\in\{\frac{1}{2},1\}$] rests on a recursive formulation of perturbation theory and the exact solution of the up to $n_{max}+2$-fold $\tau$-integrals using the kernel trick [38], see Ref. 30 for details. There we also discuss results for the static spin susceptibility, $\sim G_{ii^{\prime}}(i\nu_{m}=0)$. Note that the high expansion order $n_{max}=12$ sets our approach apart from alternative spin-diagrammatic approaches [39, 40, 41].
@@
where the termination function $\Gamma_{a,b}(s)$ is a fraction of Hermite polynomials [45] (see End Matter) and allows to take the limit $0^{+}$ in Eq. (3) exactly. The different small markers in Fig. 1(a) correspond to various linear extrapolation schemes (c.f. figure caption) to which the resulting DSF shown in various linestyles in Fig. 1(b) is practically insensitive. The so obtained DSF is in excellent agreement with DMRG [9] throughout the Brillouin zone (BZ), see Fig. 2, left column.
We now turn to $T<\infty$ $(x>0)$ where Dyn-HTE has its main advantage. Recall that it provides the expansion of moments $m_{\mathbf{k},2r}(x)$ in $x$ to order $n_{max}-2r$. Hence, for the interesting regime $x\gtrsim 1$ beyond convergence of the bare series, we apply resummation in form of Padé approximants [21] to $m_{\mathbf{k},2r}(x)$ with $r\leq 3$ and then proceed as above, see End Matter (Fig. 5) for details. Analogous to conventional HTE [46, 21] we improve convergence by changing to the auxiliary variable $u=\tanh(fx)$ with tuning parameter $f$ chosen such that different Padé approximants agree. In principle, $f$ should be as small as possible to reach large $x$ and could vary with $\mathbf{k}$ but for the chain a single $f=0.48$ works well. The 2nd and 3rd column of Fig. 2 compare the DSF at temperatures $x=2$ and $4$ from Dyn-HTE (top) to DMRG [9] (middle). Linecuts (bottom) show only minor deviations.
$S=1/2$ Heisenberg AFMs in 2D.—While magnetic long-range order is only possible at $T=0$ for these nearest-neighbor models, its precursor in the form of spin-waves can dominate the DSF already at temperatures $T\lesssim J$ [47]. We refer to the End Matter for representative Dyn-HTE results for the DSF of the square lattice model and the frustrated kagome-lattice which lacks long-range magnetic order at $T=0$ and shows broad features in the DSF.
@@
The DSF from Dyn-HTE shown in Fig. 3 (see also End Matter for a BZ-path at $x=3$) is ideally suited to shed new light on the still debated nature of this intermediate-$T$ regime. One proposed scenario starts from the $T=0$ excitation spectrum which shows a “roton-like excitation” (RLE) characterized by a dispersion minimum at the $M$-point (center of BZ edge) with gap $\Delta\simeq 0.55J$ [52, 53]. Is the intermediate-$T$ region primarily characterized by thermal excitations of these enigmatic RLEs? At $\mathbf{k}=M$, the $T$-dependence of the equal-time structure factor shows a weak maximum around $T\simeq\Delta$, see Ref. 51, but the DSF in Fig. 3(top) does not soften significantly across the intermediate $T$-range with the peak remaining around $\omega_{max}\simeq J$ [for $S(\mathbf{k}\!=\!M,\omega\!\rightarrow\!0)$, a weak maximum can however be found as $T$ decreases]. Hence, if the RLE at $\mathbf{k}=M$ is responsible for the anomalous behavior, the associated fluctuations are not simple spin-waves but must be more complex. One candidate are triangle-based chiral fluctuations [51] but the associated chiral DSF is beyond scope of this work.
An alternative scenario links the anomalous intermediate-$T$ regime to the passage through a critical fan [54] of a $T=0$ quantum phase transition (QPT) at which the $120^{\circ}$ order melts [55, 56, 50]. The QPT appears in an extended model parameter space, one candidate is the $J_{1}$-$J_{2}$-model with a QSL at $J_{2}\gtrsim 0.06J_{1}$ [57]. A hallmark of the critical-fan scenario is a temperature-frequency scaling relation for the DSF at ordering wavevector $\mathbf{k}=K$ [55] for $|\omega|\lesssim J$,
@@
where $\alpha=(2-\eta)/z$ is determined by the (unknown) critical exponents of the QPT and $\Phi(\cdot)$ is a scaling function. Interestingly, the $\mathbf{k}=K$ DSF from Dyn-HTE shown in Fig. 3(bottom) fulfills such a relation for $\alpha=1.10(2)$ (inset), but the $T$-range for which scaling occurs is relatively narrow. If scaling holds indeed and the lower limit on $T$ is not due to inaccuracies of Dyn-HTE, extensions to the $J_{1}$-$J_{2}$-model should be able to tune closer to criticality and thereby widen the $T$-range used for scaling. Other open questions pertain to the incompatible value of $\alpha\simeq 1.73(12)$ found from a similar scaling analysis of INS data from $\mathrm{KYbSe}_{2}$ with finite and slightly subcritical $J_{2}$ [58] and an extended scaling analysis taking into account deviations from the ordering wavevector, $\mathbf{k}\neq K$.
@@
S=1S\!=\!1 Heisenberg AFMs: Chain and pyrochlore lattice.—Finally, we obtain the DSF from Dyn-HTE for spin $S=1$ Heisenberg AFMs. For the chain at $x=4$, some aspects like the dispersion maxima or Haldane gap already agree well with ground-state ($x\!\rightarrow\!\infty$) DMRG results, see End Matter (Fig. 7). The $S=1$ pyrochlore material $\mathrm{NaCaNi}_{2}\mathrm{F}_{7}$ was reported to approximately realize the nearest-neighbor Heisenberg AFM [16, 59] but also possible biquadratic interactions were discussed [60]. We assume $J=$2.4\text{\,}\mathrm{meV}$\$ [60] and compare the DSF from Dyn-HTE along the $[22l]$ slice with the experimental INS result [59], see Fig. 4. We find fair agreement of the main features including the size of the “V” shape as well as the height of the dome-shape which was not captured in semiclassical studies [60, 16]. In Dyn-HTE, the maximum along the line-cut at $[220]$ appears at slightly lower $\omega$ compared to the INS data. Possible reasons are potential beyond-Heisenberg perturbations in the material and a mismatch in temperature ($x=4$ in Dyn-HTE versus $x\simeq 15$ in INS). Indeed the peak position in Dyn-HTE has a trend to higher $\omega$ with lower $T$.
Figure 4: (a) DSF for the pyrochlore lattice $S=1$ Heisenberg AFM for $\mathbf{k}=(2,2,l)$, $J=$2.4\text{\,}\mathrm{meV}$\$ at $x=J/T=4$ from Dyn-HTE (via u-Padé with $f=0.6$). (b) Experimental INS data for $\mathrm{NaCaNi}_{2}\mathrm{F}_{7}$ at $x\simeq 15$, from Ref. 59.
Conclusion.—Compared to static observables from conventional HTE, the DSF obtained via Dyn-HTE accesses a wealth of additional information on dipolar excitations. Depending on the model, the results can be trusted up to $x=J/T\simeq 4$ but advanced resummation strategies [61] might extend this range.
Looking ahead, Dyn-HTE will not only be useful to model INS data but also informs experiments with local probes like muon spin relaxation [4], inelastic electron tunneling- or noise-spectroscopy for 2D quantum magnets [62, 63, 64]. Other applications include finite-$T$ spin diffusion and the computation of quantum Fisher information as entanglement witness [6, 9]. Straightforward extensions of the Dyn-HTE formalism will consider $S>1$, models with more than one coupling constant (e.g. $J_{1}$-$J_{2}$) [65], single-ion anisotropies or magnetic fields [32]. Another worthwhile endeavor would be the extension to higher-order dynamic [66, 67] or chiral [51, 68, 40] correlators which are recently moving into focus.
@@ Acknowledgements.
Acknowledgments.—We thank Michel Ferrero, Wei Li, Matías Gonzalez, Alexander Tsirlin, Johannes Reuther, Kemp Plumb and Andreas Weichselbaum for useful discussions and the latter two for sharing DMRG data from Ref. 9 and INS data from Ref. 59, respectively. The authors acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). The authors also acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through Grant No. INST 40/575-1 FUGG (JUSTUS 2 cluster). We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Research Unit FOR 5413/1, Grant No. 465199066. B.Sch. acknowledges funding from the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. B.Sb. and B.Sch. are supported by DFG Grant No. 524270816.

### v4

@@ Abstract
Here we narrow this methodological gap by introducing a dynamic extension of the high-temperature expansion (HTE). Conventional HTE targets equal-time correlations and thermodynamic observables via graph-based expansions in $J/T$ and is developed for more than five decades [20, 21]. It benefits from a formulation in the thermodynamic limit and is oblivious to frustration, entanglement, dimensionality and spin-length [22]. Its main limitation is the model-dependent temperature range $T\gtrsim O(J/4)$. Our dynamic HTE (Dyn-HTE) shares the same benefits but applies to frequency moments of the real-frequency dynamic susceptibility from which we reconstruct the DSF [23, 24, 25]. Although these moments are considerably more complex than equal-time correlators, the high expansion orders of conventional HTE are maintained. This allows for the computation of four to seven frequency moments, depending on $T/J$, and goes beyond numerical linked-cluster expansions (NLCE) [26, 27] which computed the lowest two moments (Gauss approximation) for some particular S=1/2S\!=\!1/2 models [28].
@@
$M_{ii^{\prime},n}\!\!=\!i^{n}\partial_{t}^{n}\!\!\left\langle S_{i}^{z}(t)S_{i^{\prime}}^{z}\right\rangle\!|_{t=0}\!=\!\!\left\langle[...[[S_{i}^{z},\!H],\!H],\!...,\!H]S_{i^{\prime}}^{z}\right\rangle\!.\!\!$ (6)
@@
Dyn-HTE for moments.—We now present our main conceptual result relevant for finite $T$: The HTE of moments $m_{\mathbf{k},2r}$ is encoded in the Dyn-HTE of the Matsubara spin correlator. This is significant as the latter treats $\rho$ and imaginary-time evolution $S_{i}^{z}(\tau)=e^{H\tau}S_{i}^{z}e^{-H\tau}$ on equal footing [33] and allows for an efficient evaluation of exact HTE coefficients, see our companion article [30]. The Matsubara correlator at frequency $\nu_{m}=2\pi Tm$ ($m\in\mathbb{Z}$) and its HTE to order $n_{max}$ in x=J/Tx\!=\!J/T read
@@
Here $\mathcal{T}$ enforces $\tau$-ordering, $\Delta_{2\pi m}\equiv(1-\delta_{0,m})/(2\pi m)$ and $r_{max}=\left\lfloor n_{max}/2\right\rfloor$. The polynomials $p_{ii^{\prime}}^{(2r)}(x)$ will be linked to the HTE of the moments momentarily. They are of degree $n_{max}-2r$ with rational coefficients. Our open-source implementation [29] for $n_{max}=12$ provides $p_{ii^{\prime}}^{(2r)}(x)$ for all site-pairs $ii^{\prime}$ on arbitrary lattices $\mathcal{L}$. This flexibility hinges on the pre-evaluation of Dyn-HTE (8) on generic lattice snippets called “graphs” which are then embedded in arbitrary (and also high-dimensional and frustrated) $\mathcal{L}$ in a post-processing step as in conventional HTE [21]. Efficient evaluation of all $\sim\!10^{6}$ graphs with up to $n_{max}$ edges [currently for $S\in\{\frac{1}{2},1\}$] rests on a recursive formulation of perturbation theory and the exact solution of the $n_{max}\!+\!2$-fold $\tau$-integrals using the kernel trick [38], see Ref. 30 for details. There we also discuss results for the static spin susceptibility, $\sim G_{ii^{\prime}}(i\nu_{m}=0)$.
@@
where the termination function $\Gamma_{a,b}(s)$ is a fraction of Hermite polynomials [42] (see End Matter) and allows to take the limit $0^{+}$ in Eq. (3) exactly. The different small markers in Fig. 1(a) correspond to various linear extrapolation schemes (c.f. figure caption). The main features of the resulting DSFs shown in various linestyles in Fig. 1(b) are practically insensitive to the chosen extrapolation and is in excellent agreement with DMRG [9] throughout the Brillouin zone (BZ), see Fig. 2, left column.
We now turn to $T<\infty$ $(x>0)$ where Dyn-HTE has its main advantage. Recall that it provides the expansion of moments $m_{\mathbf{k},2r}(x)$ in $x$ to order $n_{max}-2r$. Hence, for the interesting regime $x\gtrsim 1$ beyond convergence of the bare HTE series, we apply resummation in form of Padé approximants [21] to $m_{\mathbf{k},2r}(x)$ with $r\leq 3$ and then proceed as above, see End Matter (Fig. 5) for details. Analogous to conventional HTE [43, 21] we improve convergence by changing to the auxiliary variable $u=\tanh(fx)$ with tuning parameter $f$ chosen such that different Padé approximants agree. In principle, $f$ should be as small as possible to reach large $x$ and could vary with $\mathbf{k}$ but for the chain a single $f=0.48$ works well. The 2nd and 3rd column of Fig. 2 compare the DSF at temperatures $x=2$ and $4$ from Dyn-HTE (top) to DMRG [9] (middle). Linecuts (bottom) show only minor deviations.
$S=1/2$ Heisenberg AFMs in 2D.—While magnetic long-range order is only possible at $T=0$ for nearest-neighbor models, its precursor in the form of spin-waves can dominate the DSF already at $T\lesssim J$ [44]. We refer to the End Matter for representative Dyn-HTE results for the DSF of the square lattice model and the frustrated kagome lattice which lacks long-range magnetic order at $T=0$ and shows broad features in the DSF.
@@
The DSF from Dyn-HTE shown in Fig. 3 (see also End Matter for a BZ-path at $x=3$) is ideally suited to shed new light on the still debated nature of this intermediate-$T$ regime. One proposed scenario starts from the $T=0$ excitation spectrum which shows a “roton-like excitation” (RLE) characterized by a dispersion minimum at the $M$-point (center of BZ edge) with gap $\Delta\simeq 0.55J$ [49, 50]. Is the intermediate-$T$ region primarily characterized by thermal excitations of these enigmatic RLEs? At $\mathbf{k}=M$, the $T$-dependence of the equal-time structure factor shows a weak maximum around $T\simeq\Delta$, see Ref. 48, but the DSF in Fig. 3(top) does not soften significantly across the intermediate $T$-range with the peak remaining around $\omega_{max}\simeq J$ [for $S(\mathbf{k}\!=\!M,\omega\!\rightarrow\!0)$, a weak maximum can however be found as $T$ decreases]. Hence, if the RLE at $\mathbf{k}=M$ is responsible for the anomalous behavior, the associated fluctuations are not simple spin-waves but must be more complex. One candidate are triangle-based chiral fluctuations [48] but the associated chiral DSF is beyond scope of this work.
An alternative scenario links the anomalous intermediate-$T$ regime to passage through a critical fan [51] of a $T=0$ quantum phase transition (QPT) at which the $120^{\circ}$ order melts [52, 53, 47]. A hallmark of this scenario is a temperature-frequency scaling relation for the DSF at ordering wavevector $\mathbf{k}\!=\!K$ for $|\omega|\lesssim J$,
@@
where $\alpha=(2-\eta)/z$ is determined by the critical exponents of the QPT and $\Phi(\cdot)$ is a scaling function. Interestingly, the $\mathbf{k}=K$ DSF from Dyn-HTE shown in Fig. 3(bottom) fulfills such a relation for $\alpha=1.10(2)$ (inset), but the $T$-range for which scaling occurs is relatively narrow. One candidate for the QPT in question is the transition to a Dirac QSL in the $J_{1}$-$J_{2}$-model at $J_{2}\simeq 0.06J_{1}$ [54]. Although the critical exponents are unknown, Lorentz invariance would imply $z=1$ and $\eta=O(1)$ has been found in the potentially related Gross-Neveu Heisenberg model via large-$N_{f}$ or $\epsilon=4-d$ expansion [55, 56]. Our value of $\alpha$ is thus not implausible and can inform further efforts on the field-theory side, although the scaling analysis should be repeated with Dyn-HTE data for the $J_{1}$-$J_{2}$-model. Other open questions pertain to the incompatible value of $\alpha\simeq 1.73(12)$ found from a similar scaling analysis of INS data from $\mathrm{KYbSe}_{2}$ with finite and slightly subcritical $J_{2}$ [57] and an extended scaling analysis taking into account deviations from the ordering wavevector, $\mathbf{k}\neq K$.
@@
S=1S\!=\!1 Heisenberg AFMs: Chain and pyrochlore lattice.—Finally, we obtain the DSF from Dyn-HTE for spin $S=1$ Heisenberg AFMs. For the chain at $x=4$, some aspects like the dispersion maxima or Haldane gap already agree well with ground-state ($x\!\rightarrow\!\infty$) DMRG results, see End Matter (Fig. 7). The $S=1$ pyrochlore material $\mathrm{NaCaNi}_{2}\mathrm{F}_{7}$ was reported to approximately realize the nearest-neighbor Heisenberg AFM [16, 58] with $J=$2.4\text{\,}\mathrm{meV}$\$ [59]. In Fig. 4, we compare the DSF from Dyn-HTE along the $[22l]$ momentum with experimental INS results [58]. We find fair agreement of the main features including the size of the “V” shape as well as the height of the dome-shape which was not captured in semiclassical studies [59, 16]. In Dyn-HTE, the maximum along the line-cut at $[220]$ appears at slightly lower $\omega$ compared to INS data. Possible reasons are potential beyond-Heisenberg perturbations in the material and a mismatch in temperature ($x=4$ in Dyn-HTE versus $x\simeq 15$ in INS). Indeed the peak position in Dyn-HTE has a trend to higher $\omega$ with lower $T$.
Figure 4: (a) DSF for the pyrochlore lattice $S=1$ Heisenberg AFM for $\mathbf{k}=(2,2,l)$, $J=$2.4\text{\,}\mathrm{meV}$\$ at $x=J/T=4$ from Dyn-HTE (via u-Padé with $f=0.6$). (b) Experimental INS data for $\mathrm{NaCaNi}_{2}\mathrm{F}_{7}$ at $x\simeq 15$, from Ref. 58.
Conclusion.—Compared to static observables of conventional HTE, the DSF obtained via Dyn-HTE accesses a wealth of additional information on dipolar excitations down to moderately-low $T\gtrsim J/4$. On a technical level we achieve this by exploiting a graph-based approach for high-order moment expansions and fine momentum resolution, combined with continued-fraction reconstruction, judiciously chosen resummation schemes all packaged in a single algorithmic recipe [29]. This combination sets our method apart from previous (spin-)diagrammatic approaches in a qualitative manner [60, 61, 62, 63, 64, 35].
Looking ahead, Dyn-HTE will not only be useful to model INS data but also informs experiments with local probes like muon spin relaxation [4], inelastic electron tunneling- or noise-spectroscopy for 2D quantum magnets [65, 66, 67]. Other applications include finite-$T$ spin diffusion and the computation of quantum Fisher information as entanglement witness [6, 9]. Straightforward extensions of the Dyn-HTE formalism will consider advanced resummation strategies [68], $S>1$, models with multiple couplings (e.g. $J_{1}$-$J_{2}$) [69], single-ion anisotropies or magnetic fields [32]. Another worthwhile extension would be higher-order dynamic [70, 71] or chiral [48, 72, 62] correlators moving into focus recently.
@@ Acknowledgements.
Acknowledgments.—We thank Michel Ferrero, Matías Gonzalez, Lukas Janssen, Wei Li, Johannes Reuther, Alexander Tsirlin, Andreas Weichselbaum, and Kemp Plumb for useful discussions and the latter two for sharing DMRG data from Ref. 9 and INS data from Ref. 58, respectively. The authors acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). The authors also acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through Grant No. INST 40/575-1 FUGG (JUSTUS 2 cluster). We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Research Unit FOR 5413/1, Grant No. 465199066. B.Sch. acknowledges funding from the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. B.Sb. and B.Sch. are supported by DFG Grant No. 524270816.